How many mL of a 6.0M HCl solution would be required to react with 0.250 of excess Mg after reaction with the Cu^2+
Answer #1

Molar mass of Mg = 24.3 g/mol

Number of moles of Mg = (Mass)/(Molar Mass) = 0.250/24.3 = 0.010288 mol

One mole of Mg reacts with two moles of HCl

Moles of HCl needed = 0.010288 * 2 = 0.020576 mol

Moles of HCl = Volume of solution( in L) * Molarity (M)

0.020576 = x * 6

x = 0.003429 L or 3.43 ml

Hence the volume of HCl needed will be 3.43 ml

Note - Post any doubts/queries in comments section.